# thanks: $0$-0002-5209-872X

### v2

@@ I Introduction
The best way to probe CP violation in the leptonic sector is by an appearance measurement of an oscillation maximum Itow et al. (2001); Marciano (2001); Dick et al. (1999); Cervera et al. (2000); Autiero et al. (2004); Nunokawa et al. (2008); Coloma et al. (2013); Denton et al. (2022). That is, the detection of one flavor of neutrino in a source of neutrino that is predominantly a differently flavor at a baseline and energy that corresponds to one of the $\Delta m^{2}_{ij}$’s. To date only NOvA Adamson et al. (2016) and T2K Abe et al. (2014) have strong evidence for the detection of appearance by detecting electron neutrinos in predominantly muon neutrino sources, but do not yet significantly probe CP violation past the $2\sigma$ level Acero et al. (2022a); Ramírez et al. (2023). Atmospheric neutrinos, which are mostly muon neutrinos also have some evidence for appearance Li et al. (2018); Jiang et al. (2019); Aartsen et al. (2019). Since there is significant electron neutrino contribution at the source, the appearance information is thus somewhat scrambled.
@@
We will show how it is also possible to probe CP violation, via neutrino disappearance measurements only, in three different ways: counting parameters using a specific parameterization, direct analytic calculation in a parameterization independent framework, and numerical computation. With good measurements of the disappearance of two different flavors it is possible to determine a total of four independent parameters in the PMNS matrix. Since the mixing matrix is only described by four parameters [^1] Two additional parameters, the so-called Majorana phases, may also be physical depending on the nature of neutrinos, but their impact in neutrino oscillation experiments is suppressed by $(m_{\nu}/E_{\nu})^{2}<10^{-14}$ or smaller in oscillations and thus can be safely ignored. it is possible to learn about CP violation while only measuring CP conserving channels. The key physics effect that makes this possible is unitarity [^2] We note that unitarity is used for most experiment’s extraction of one or more of the mixing parameters. For example, medium-baseline reactor neutrino experiments at $L\simeq 1$ km actually determine $4|U_{e3}|^{2}(|U_{e1}|^{2}+|U_{e2}|^{2})$ but by unitarity this is equal to $4|U_{e3}|^{2}(1-|U_{e3}|^{2})$ which can be expressed as $\sin^{2}2\theta_{13}$. and thus if there is new physics in the neutrino sector this story may get more complicated. It is also possible to relate the amount of CP violation directly to the measured parameters from disappearance. Given the sizable expected improvements in disappearance measurements in the $\nu_{e}$ channel with the Jiangmen Underground Neutrino Observatory (JUNO) and the $\nu_{\mu}$ channel with the Deep Underground Neutrino Experiment (DUNE) and Hyper-Kamiokande (HK), such a study is quite timely. Moreover, disappearance has a different dependence on the oscillation parameters as well as different (and often cleaner) systematics than appearance measurements since the neutrinos at the near and far detectors are the same flavor which means that this can be a valuable cross check of CP violation probes in the appearance channel. See also Farzan and Smirnov (2002) for a very early study discussing some related oscillation physics and Rout et al. (2020) for some numerical studies.
@@ II Conventional CP violation picture
New physics such as sterile neutrinos Acero et al. (2022b), non-standard neutrino interactions Wolfenstein (1978); Pro (2019); Argüelles et al. (2023), or unitarity violation Martinez-Soler and Minakata (2020) could also modify this picture in non-trivial ways by making fully CP conserving scenarios appear CP violating or other such nightmare scenarios. It may be possible to avoid these scenarios via a combination of experiments at different baselines and energies, see e.g. de Gouvêa and Kelly (2016); Denton et al. (2021); Chatterjee and Palazzo (2021).
@@ III.1 An understanding via parameter counting
While it is not directly possible to determine if nature prefers neutrinos or antineutrinos via disappearance measurements alone, it is possible to determine if nature treats neutrinos and antineutrinos the same or differently via measurements of these CP conserving disappearance channels [^5] It has been pointed out in Luo and Xing (2023) that the CP conserving parts of CP violating channels could also provide some amount of information about CP violation.. That is, disappearance measurements cannot provide information on $\operatorname{sign}(\sin\delta)$ or equivalently on $\operatorname{sign}J$, but can constrain $\cos\delta$ and thus potentially rule out CP conserving values of $|\cos\delta|=1$.
@@
To understand how one can determine if CP is conserved or not, we focus on the four parameters that describe the mixing matrix. We begin by examining the PMNS mixing matrix in the usual parameterization [^6] While we are here working in a parameterization dependent framework, the ability to discover CP violation is not artificially enhanced by this and CP violation can be discovered in any unitary parameterization of the mixing matrix because CP violation is a physical effect, see eq. 6 below. Tanabashi et al. (2018); Denton and Pestes (2021):
@@
Since disappearance measurements only constrain absolute values of elements of the PMNS matrix, we notice that the measurements of the first row, the $\nu_{e}$ row, provide no information about $\delta$. It would seem that measurements of either the $\nu_{\mu}$ or $\nu_{\tau}$ row would provide information about $\delta$, specifically $\cos\delta$, implying that $\nu_{e}$’s are somehow special and different from the other two flavors [^7] This apparent difference has been investigated in the literature, see e.g. Kimura et al. (2006); Akhmedov et al. (2008), but the importance of two simultaneous disappearance measurements has not been realized.. In reality, any one row (and any one column) can be made to be “simple”: only a product of sines, cosines, and $e^{\pm i\delta}$, see e.g. Denton and Pestes (2021). The remaining four elements must always be “complicated”: the sum or difference of the products of such terms, one of which always contains $e^{\pm i\delta}$. This provides one means of understanding why two separate disappearance measurements are required to probe CP violation. That is, if, for example, we had an excellent measurement of $\nu_{\mu}$ disappearance but not $\nu_{e}$ or $\nu_{\tau}$ disappearance, since we could choose to make the $\nu_{\mu}$ row simple then therefore we cannot learn anything about CP violation.
@@
Since the appearance channel essentially constrains $\sin\delta$ (see eq. 2) while the disappearance channel constrains $\cos\delta$, these two measurements provide key complementary information. In fact, there will be sign degeneracies in many regions of parameter space of $\delta$ with either only appearance or disappearance. Moreover, the precision on $\delta$ near $\pi/2$ or $23\pi/$ 2 is determined by the sensitivity to $\cos\delta$ which comes from this combination of disappearance measurements making this disappearance based measurement crucial for determining the exact value of $\delta$ if we are near $|\sin\delta|=1$ as some data Ramírez et al. (2023) may be indicating.
@@ III.2 A direct analytic calculation
We now leverage approximation techniques to theoretically investigation of the size of the effect.
@@ III.3 Analytic approximation
where the $\cos\delta$ dependence is numerically $\approx-0.0005\cos\delta$ at $E_{\max}=1.3$ GeV, the energy of the first non-trivial maximum for DUNE. Third, we include the correction due to the matter effect which significantly changes this. The matter effect has almost no impact on $\theta_{23}$ or $\delta$ below the atmospheric resonance at $E\simeq 11$ GeV Denton et al. (2016); Xing et al. (2018). In addition, while $\theta_{13}$, $\Delta m^{2}_{31}$, and $\Delta m^{2}_{32}$ do evolve somewhat in matter, they change $\lesssim 10\%$ from their vacuum values and the effect can be safely ignored for the $\Delta_{21}$ term. The solar parameters, $\theta_{12}$ and $\Delta m^{2}_{21}$, on the other hand, evolve considerably in matter at these energies. To a sufficient approximation, the matter correction factor for the solar parameters is Denton et al. (2016); Denton and Parke (2018)
@@
where $a=2\sqrt{2}G_{F}N_{e}E$ is the contribution from the matter effect. At $E_{\max}=1.3$ GeV, $\mathcal{S}_{\odot}=3.4$ and provides an excellent approximation to the rescaling of $\Delta m^{2}_{21}$ in matter. We note that past the solar resonance at $E=0.13$ GeV, $\theta_{12}>\pi/4$ and thus $\cos 2\theta_{12}<0$ flipping the sign on the $\cos\delta$ dependence. We can approximate the solar mixing angle in matter by
@@
For antineutrinos the story is somewhat different. The value of $\cos\delta$ does not change as $\delta\to-\delta$, so it is the same as for neutrinos. In matter we note that while $\cos 2\widehat{\theta_{12}}<0$ for neutrinos for DUNE and HK, it remains positive for antineutrinos, as in vacuum. Thus the impact on the oscillation maxima for antineutrinos is the same in matter as in vacuum and the probability is comparatively large for $\cos\delta=-1$ and small for $\cos\delta=+1$. Due to the lower statistics in $\bar{\nu}_{\mu}$ mode however, this channel will not contribute as much as the neutrino channel to the total significance for probing $\cos\delta$ and CP violation.
@@ IV Estimated experimental sensitivities
To numerically quantify the magnitude of the effect given realistic experimental details, we simulate $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ disappearance in DUNE using DUNE’s simulation files Abi et al. (2021); Huber et al. (2005). We consider 40 kt fiducial volume and 6.5 years in each neutrino and antineutrino mode with 1.2 MW and 56% beam uptime. We consider priors on the five oscillation parameters other than $\delta$ from one of
@@
    our current knowledge of the oscillation parameters Esteban et al. (2020). Since both DUNE and HK will provide better measurements of $\theta_{23}$ and $\Delta m^{2}_{31}$ than existing data, this is equivalent to using Daya Bay, RENO, KamLAND, and Solar data (all disappearance experiments) to constrain $\theta_{13}$, $\theta_{12}$, and $\Delta m^{2}_{21}$,
@@
    the expected improvement on $\theta_{12}$, $\Delta m^{2}_{21}$, and $\Delta m^{2}_{31}$ from the inclusion of 6 years of JUNO’s $\bar{\nu}_{e}$ disappearance data Abusleme et al. (2022), and
@@
Figure 2: The expected sensitivity to $\cos\delta$ using only $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ disappearance from DUNE along with external priors on the other five oscillation parameters from: the current precision, the expected improvement with JUNO, or if the other oscillation are known perfectly.
We have confirmed that there is information about $\cos\delta$ in each of neutrino and antineutrino modes individually, although neutrino mode contributes more to the information due to higher statistics from the larger cross section and lower wrong-sign lepton rates. In addition, as suggested by the theory discussion above, $\cos\delta$ can also be determined if DUNE was performed in vacuum, although the results would be modified. Additional numerical results for DUNE as well as HK can be found in the appendix including the impact of runtime.
@@
One could also consider $\nu_{\mu}$ disappearance with atmospheric neutrinos at HK Abe et al. (2018), IceCube Abbasi et al. (2023), KM3NeT Adrian-Martinez et al. (2016), or JUNO Abusleme et al. (2021); Suliga and Beacom (2023), however the expected sensitivity is likely less than that presented here and depends strongly on systematics, see e.g. Argüelles et al. (2022) for a discussion including both disappearance and appearance in atmospherics. Nonetheless, due to the different systematics and timelines it may be useful to consider a fit including atmospheric neutrinos alongside state-of-the-art $\nu_{e}$ disappearance measurements.
In principle, one could probe $\cos\delta$ with existing disappearance data. The current status of the data is that the best $\nu_{\mu}$ disappearance measurements come from NOvA Acero et al. (2022a) and T2K Ramírez et al. (2023) and the best $\nu_{e}$ disappearance measurements come from Daya Bay Adey et al. (2018) and KamLAND Gando et al. (2013). The $\nu_{e}$ disappearance data is described by the “current” curves in fig. 2 as well as those in the appendix which show that even DUNE or HK can only provide at most $\sim 1.4\sigma$ sensitivity to $\cos\delta$; with existing NOvA and T2K data there would not be any significant $\cos\delta$ information at all.
@@ Acknowledgements.
Appendix A APPENDIX
Appendix B CP violation discovery sensitivity
We now include, for completeness, several additional numerical results including both HK and DUNE.
In fig. 3, we quantify the expected sensitivity to discover CP violation (that is, ruling out $|\cos\delta|=1$) as a function of the true value of $\delta$ for DUNE (left) and HK (right). We also consider different choices of external measurements of the other oscillation parameters as in fig. 2. The different colors correspond to including both appearance and disappearance (the standard DUNE analysis), only appearance, and only disappearance. The different line styles correspond to external pulls from the current knowledge of the oscillation parameters, the expected improvements with JUNO, and hypothetical perfect knowledge. The blue dotted curve (both channels and current knowledge of the oscillation parameters) agrees with DUNE’s curve very well.

### v3

@@ I Introduction
The best way to probe CP violation in the leptonic sector is by an appearance measurement of an oscillation maximum Itow et al. (2001); Marciano (2001); Dick et al. (1999); Cervera et al. (2000); Autiero et al. (2004); Nunokawa et al. (2008); Coloma et al. (2013); Denton et al. (2022). That is, the detection of one flavor of neutrino in a source of neutrino that is predominantly a differently flavor at a baseline and energy that corresponds to one of the $\Delta m^{2}_{ij}$’s. To date only NOvA Adamson et al. (2016) and T2K Abe et al. (2014) have strong evidence for the detection of appearance by detecting electron neutrinos in predominantly muon neutrino sources, but do not yet significantly probe CP violation past the $2\sigma$ level Acero et al. (2022a); Abe et al. (2023). Atmospheric neutrinos, which are mostly muon neutrinos also have some evidence for appearance Li et al. (2018); Jiang et al. (2019); Aartsen et al. (2019). Since there is significant electron neutrino contribution at the source, the appearance information is thus somewhat scrambled.
@@
We will show how it is also possible to probe CP violation, via neutrino disappearance measurements only, in three different ways: counting parameters using a specific parameterization, direct analytic calculation in a parameterization independent framework, and numerical computation. With good measurements of the disappearance of two different flavors it is possible to determine a total of four independent parameters in the PMNS matrix. Since the mixing matrix is only described by four parameters it is possible to learn about CP violation while only measuring CP conserving channels. The key physics effect that makes this possible is unitarity [^1] We note that unitarity is used for most experiment’s extraction of one or more of the mixing parameters. For example, medium-baseline reactor neutrino experiments at $L\simeq 1$ km actually determine $4|U_{e3}|^{2}(|U_{e1}|^{2}+|U_{e2}|^{2})$ but by unitarity this is equal to $4|U_{e3}|^{2}(1-|U_{e3}|^{2})$ which can be expressed as $\sin^{2}2\theta_{13}$. and thus if there is new physics in the neutrino sector this story may get more complicated. It is also possible to relate the amount of CP violation directly to the measured parameters from disappearance. Given the sizable expected improvements in disappearance measurements in the $\nu_{e}$ channel with the Jiangmen Underground Neutrino Observatory (JUNO) and the $\nu_{\mu}$ channel with the Deep Underground Neutrino Experiment (DUNE) and Hyper-Kamiokande (HK), such a study is quite timely. Moreover, disappearance has a different dependence on the oscillation parameters as well as different (and often cleaner) systematics than appearance measurements since the neutrinos at the near and far detectors are the same flavor which means that this can be a valuable cross check of CP violation probes in the appearance channel. See also Farzan and Smirnov (2002) for a very early study discussing some related oscillation physics and Rout et al. (2020) for some numerical studies.
@@ II Conventional CP violation picture
New physics such as sterile neutrinos Acero et al. (2022b), non-standard neutrino interactions Wolfenstein (1978); Pro (2019); Argüelles et al. (2023a), or unitarity violation Martinez-Soler and Minakata (2020) could also modify this picture in non-trivial ways by making fully CP conserving scenarios appear CP violating or other such nightmare scenarios. It may be possible to avoid these scenarios via a combination of experiments at different baselines and energies; see e.g. de Gouvêa and Kelly (2016); Denton et al. (2021); Chatterjee and Palazzo (2021).
@@ III.1 An understanding via parameter counting
While it is not directly possible to determine if nature prefers neutrinos or antineutrinos via disappearance measurements alone, it is possible to determine if nature treats neutrinos and antineutrinos the same or differently via measurements of these CP conserving disappearance channels [^4] It has been pointed out in Luo and Xing (2023) that the CP conserving parts of appearance channels could also provide some amount of information about CP violation.. That is, disappearance measurements cannot provide information on $\operatorname{sign}(\sin\delta)$ or, equivalently, on $\operatorname{sign}J$, but can constrain $\cos\delta$ and thus potentially rule out CP conserving values of $|\cos\delta|=1$.
@@
To understand how one can determine if CP is conserved or not, we focus on the four parameters that describe the mixing matrix [^5] Two additional parameters, the so-called Majorana phases, may also be physical depending on the nature of neutrinos, but their impact in neutrino oscillation experiments is suppressed by $(m_{\nu}/E_{\nu})^{2}<10^{-14}$ or smaller in oscillations and thus can be safely ignored.. We begin by examining the PMNS mixing matrix $U$ in the usual parameterization [^6] While we are here working in a parameterization dependent framework, the ability to discover CP violation is not artificially enhanced by this and CP violation can be discovered in any unitary parameterization of the mixing matrix because CP violation is a physical effect, see eq. 6 below. Tanabashi et al. (2018); Denton and Pestes (2021):
@@
Since disappearance measurements only constrain absolute values of elements of the PMNS matrix, we notice that the measurements of the first row, the $\nu_{e}$ row, provide no information about $\delta$. It would seem that measurements of either the $\nu_{\mu}$ or $\nu_{\tau}$ row would provide information about $\delta$, specifically $\cos\delta$, implying that $\nu_{e}$’s are somehow special and different from the other two flavors [^7] This apparent difference has been investigated in the literature, see e.g. Kimura et al. (2006); Akhmedov et al. (2008), but the importance of two simultaneous disappearance measurements has not been realized.. In reality, any one row (and any one column) can be made to be “simple”: only a product of sines, cosines, and $e^{\pm i\delta}$, see, e.g., Denton and Pestes (2021). The remaining four elements must always be “complicated”: the sum or difference of the products of such terms, one of which always contains $e^{\pm i\delta}$. This provides one means of understanding why two separate disappearance measurements are required to probe CP violation. That is, if, for example, we had an excellent measurement of $\nu_{\mu}$ disappearance but not $\nu_{e}$ or $\nu_{\tau}$ disappearance, since we could choose to make the $\nu_{\mu}$ row simple then therefore we cannot learn anything about CP violation.
@@
Since the appearance channel essentially constrains $\sin\delta$ (see eq. 2) while the disappearance channel constrains $\cos\delta$, these two measurements provide key complementary information. In fact, there will be sign degeneracies in many regions of parameter space of $\delta$ with either only appearance or disappearance. Moreover, the precision on $\delta$ near $\pi/2$ or $23\pi/$ 2 is determined by the sensitivity to $\cos\delta$ which comes from this combination of disappearance measurements making this disappearance based measurement crucial for determining the exact value of $\delta$ if we are near $|\sin\delta|=1$ as some data Abe et al. (2023) may be indicating.
@@ III.2 A direct analytic calculation
We now leverage approximation techniques to theoretically investigate the size of the effect.
@@ III.3 Analytic approximation
where the $\cos\delta$ dependence is numerically $\approx-0.0005\cos\delta$ at $E_{\max}=1.3$ GeV, the energy of the first non-trivial maximum for DUNE. Thus we would expect that the maxima will be shifted lower for $\cos\delta=1$ and higher for $\cos\delta=-1$. Third, we include the correction due to the matter effect which significantly changes this. The matter effect has almost no impact on $\theta_{23}$ or $\delta$ below the atmospheric resonance at $E\simeq 11$ GeV Denton et al. (2016); Xing et al. (2018). In addition, while $\theta_{13}$, $\Delta m^{2}_{31}$, and $\Delta m^{2}_{32}$ do evolve somewhat in matter, they change $\lesssim 10\%$ from their vacuum values and the effect can be safely ignored for the $\Delta_{21}$ term. The solar parameters, $\theta_{12}$ and $\Delta m^{2}_{21}$, on the other hand, evolve considerably in matter at these energies. To a sufficient approximation, the matter correction factor for the solar parameters is Denton et al. (2016); Denton and Parke (2018)
@@
where $a=2\sqrt{2}G_{F}N_{e}E$ is the contribution from the matter effect. At $E_{\max}=1.3$ GeV, $\mathcal{S}_{\odot}=3.4$ and provides an excellent approximation to the rescaling of $\Delta m^{2}_{21}$ in matter. See the appendix for more discussion of the solar corrections including higher order terms. We note that past the solar resonance at $E=0.13$ GeV, $\theta_{12}>\pi/4$ and thus $\cos 2\theta_{12}<0$ flipping the sign on the $\cos\delta$ dependence. We can approximate the solar mixing angle in matter by
@@
For antineutrinos the story is somewhat different. The value of $\cos\delta$ does not change as $\delta\to-\delta$, so it is the same as for neutrinos. In matter we note that while $\cos 2\widehat{\theta_{12}}<0$ for neutrinos for DUNE and HK, it remains positive for antineutrinos, as in vacuum. Thus the impact on the oscillation maxima for antineutrinos is the same in matter as in vacuum and the probability is comparatively large for $\cos\delta=-1$ and small for $\cos\delta=+1$. Because of the lower statistics in $\bar{\nu}_{\mu}$ mode however, this channel will not contribute as much as the neutrino channel to the total significance for probing $\cos\delta$ and CP violation.
@@ IV Estimated experimental sensitivities
To numerically quantify the magnitude of the effect given realistic experimental details, we simulate $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ disappearance in DUNE using DUNE’s simulation files Abi et al. (2021); Huber et al. (2005). We consider 40 kt fiducial volume and 6.5 years in each neutrino and antineutrino mode with 1.2 MW and 56% beam uptime. We consider priors on the five oscillation parameters other than $\delta$ from one of the following:
@@
    Our current knowledge of the oscillation parameters Esteban et al. (2020). Since both DUNE and HK will provide better measurements of $\theta_{23}$ and $\Delta m^{2}_{31}$ than existing data, this is equivalent to using Daya Bay, RENO, KamLAND, and Solar data (all disappearance experiments) to constrain $\theta_{13}$, $\theta_{12}$, and $\Delta m^{2}_{21}$.
@@
    The expected improvement on $\theta_{12}$, $\Delta m^{2}_{21}$, and $\Delta m^{2}_{31}$ from the inclusion of 6 years of JUNO’s $\bar{\nu}_{e}$ disappearance data Abusleme et al. (2022).
@@
Figure 2: The expected sensitivity to $\cos\delta$ using only $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ disappearance from DUNE along with external priors on the other five oscillation parameters from the current precision, the expected improvement with JUNO, or if the other oscillation parameters are known perfectly.
We have confirmed that there is information about $\cos\delta$ in each of neutrino and antineutrino modes individually, although neutrino mode contributes more to the information due to higher statistics from the larger cross section and lower wrong-sign lepton rates. In addition, as suggested by the theory discussion above, $\cos\delta$ can also be determined if DUNE was performed in vacuum, although the results would be modified. Numerous additional numerical results for DUNE as well as HK can be found in the appendix including the impact of runtime.
@@
One could also consider $\nu_{\mu}$ disappearance with atmospheric neutrinos at HK Abe et al. (2018), IceCube Abbasi et al. (2023), KM3NeT Adrian-Martinez et al. (2016), or JUNO Abusleme et al. (2021); Suliga and Beacom (2023), however the expected sensitivity is likely less than that presented here and depends strongly on systematics, see e.g. Argüelles et al. (2023b) for a discussion including both disappearance and appearance in atmospherics. Nonetheless, due to the different systematics and timelines it may be useful to consider a fit including atmospheric neutrinos alongside state-of-the-art $\nu_{e}$ disappearance measurements.
In principle, one could probe $\cos\delta$ with existing disappearance data. The current status of the data is that the best $\nu_{\mu}$ disappearance measurements come from NOvA Acero et al. (2022a) and T2K Abe et al. (2023) and the best $\nu_{e}$ disappearance measurements come from Daya Bay Adey et al. (2018) and KamLAND Gando et al. (2013). The $\nu_{e}$ disappearance data is described by the “current” curves in fig. 2 as well as those in the appendix which show that even DUNE or HK can only provide at most $\sim 1.4\sigma$ sensitivity to $\cos\delta$; with existing NOvA and T2K data there would not be any significant $\cos\delta$ information at all.
@@ Acknowledgements.
[…]
oneΔ
VI Appendix
Appendix A CP violation discovery sensitivity
We now include, for completeness, several additional numerical results including both HK and DUNE. In fig. 3, we quantify the expected sensitivity to discover CP violation (that is, ruling out $|\cos\delta|=1$) as a function of the true value of $\delta$ for DUNE (left) and HK (right). We also consider different choices of external measurements of the other oscillation parameters as in fig. 2. The different colors correspond to including both appearance and disappearance (the standard DUNE analysis), only appearance, and only disappearance. The different line styles correspond to external pulls from the current knowledge of the oscillation parameters, the expected improvements with JUNO, and hypothetical perfect knowledge. The blue dotted curve (both channels and current knowledge of the oscillation parameters) agrees with DUNE’s curve very well.